OpenStudy (anonymous):

What is the simplified form of each expression? 7^4

OpenStudy (anonymous):

7^4=7*7*7*7

OpenStudy (anonymous):



OpenStudy (anonymous):

So is that the same answer

OpenStudy (anonymous):

i think that your answer is 2,401... is that an option?

OpenStudy (anonymous):

Why simply give him/her the answer Kelly226?

OpenStudy (anonymous):

sorry...

OpenStudy (anonymous):

Ms Kelly right but can you explain

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thanks, just new to this math

OpenStudy (anonymous):

7^4 is the same as 7*7*7*7 which is the same as 49*49

OpenStudy (anonymous):

Oh ok. :) I see now

OpenStudy (anonymous):

a^b = a "times itself b times"

OpenStudy (anonymous):

Oh yeah just went back to the book Thanks!!!